Octane has a vapor pressure of 40.0 torr at 45.1°C. What would be its vapor pressure at 104.0°C? Heat of vaporization () of octane 39 kJ/mol.

200. torr
300. torr
400. torr
500. torr
Answer #1

-to Ti= u5.1% = 45.1 +273 = 318.1K Tz = 104.0°C = 104 +213 = 371K in 8 = Alvar (t =) In Pie - 39 KSynd t en R = 4490.9 (4.9*1